A grasshopper jumps a horizontal distance of 1.40 m from rest, with an initial velocity at a 50.0° angle with respect to the horizontal. (a) Find the initial speed of the grasshopper. m/s (b) Find the maximum height reached. m
Answer #1

given

x=1.4m

\Theta =50

g=9.8m/sec2

a)

u=\sqrt{\frac{xg}{sin2\Theta }}

u=\sqrt{\frac{1.4\times 9.8}{sin(2\times 50)}}

u=3.73m/sec

in three significant figure

---------------------------------------------------------

b)the ver

tical component of velocity

v=usin50=3.73m/sec\times sin50=2.86m/sec

v=2.86m/sec(in three significant figure)

so the maximum height

h=\frac{v^{2}}{2g}

h=\frac{2.86^{2}}{2\times 9.8}

h=0.417m

in three significant figure